# Sound Unblending: Exploring Sound Manipulations for Accessible Mixed-Reality Awareness

### v1

@@ 5.1. Participants
Through word-of-mouth and public recruitment posts, we recruited 18 BVI participants (10 M and 8 F), aged 20 to 41 (mean=29.0), with a deep experience of using sounds in their lives. All had professional O&M training and various sound-related experiences in other domains, like playing instruments, being voice actors, participating in orchestras, and broadcasting. Twelve were blind since birth, while six lost their vision later in life (Table 1). We refer to our BVI participants as B1-B18 in the following sections.
Table 1. Participant demographics information.
@@
B16 29 Female Blind, since birth. Light perception. O&M, Instrument, Making AD
@@ 5.2. Simulated Scenarios and Sound Manipulations
To fully explore the space of MR, we present three scenarios with different focuses on sound awareness: RW-Focused, VR-Focused, and Fully-Mixed. We simulated the three scenarios in Unity and used its built-in audio functions to prototype the six manipulators. In each scenario, there were four types of sounds, and participants were asked to complete sound identification tasks by pressing down specific keys on the keyboard upon hearing corresponding sounds.
@@
Simulated RW-Focused Scenario. (a) The user’s avatar in Unity wears a headphone and holds a white cane. (b) The user sets the sound output by placing audio directions on the left, ringtone on the right, and (c) navigates on the street. (d) Several crowds, vendors, and passing cars along the street generate ambient noises, as well as (e) construction sites with drilling noises. (f) While walking, the user might come across random manholes, causing the white cane’s sound to change upon contact.
@@
  • Ambience Builder: applies half acoustic transparency (Figure 1d.1).
  • Feature Shifter: increases the volume of white cane sounds (Figure 1d.2).
  • Spatializer: places navigation instructions on the left and ringtones on the right (Figure 1d.5).
  • Stylizer: applies low pass filters to drilling noises to be audibly comfortable (Figure 1d.3).
  • Earcon Generator: appends a short earcon when getting close to a construction site (Figure 1d.3).
  • Prioritizer: prioritizes the four sound types by volume: white cane on the manhole, drilling, navigation instructions, and ringtone. Additionally, if RW sounds occur, virtual sounds are delayed until after the RW sounds end (see Figure 1d.4).
@@
Simulated VR-Focused Scenario. (a) The user’s avatar in Unity wears headphones, sits and works at the help desk, listens to an audio handbook, and (b) occasionally receives voice notes from a supervisor. (c) In the environment, there are background noises when people walk around, talk to each other, and open/close the sliding door. (d) People sometimes knock on the desk to get the user’s attention. (e) The speaker plays occasional public announcements on the front wall.
@@
  • Ambience Builder: applies full noise cancellation by default and dynamically switches to half acoustic transparency during public announcements.
  • Feature Shifter: increases the volume of knocking and public announcements.
  • Spatializer: places audio handbook on the left and voice notes on the right.
  • Prioritizer: prioritizes the four sound types by volume: knocking, public announcements, audio handbook, and voice notes. Additionally, if RW sounds occur, virtual sounds are delayed until after the RW sounds end.
@@
Simulated Fully-Mixed Scenario. (a) The user’s avatar sits at the dining table and wears headphones to consume the voice of virtual speakers and virtual broadcasts. (b) There are physical and remote virtual speakers on the front stage, (c) where they may stand and speak up at the same time in a panel discussion, (d) similar to the physical and virtual attendees around the table. (e) Waitstaff sometimes comes to clean the table, generating dish clinking sounds.
@@
  • Ambience Builder: applies half acoustic transparency.
  • Spatializer: places virtual broadcasts on the right.
  • Stylizer: applies a heavier old-school telephone effect to virtual voices.
  • Earcon Generator: appends two separate earcons for table cleaning and virtual broadcasts.
  • Prioritizer: prioritizes the four sound types using volume in the order of virtual/real people’s voices, table cleaning, and virtual broadcasts. Furthermore, if RW sounds are happening, virtual sounds will be delayed after RW sounds end.
5.3. Technical Details in Unity Implementation
In each trial, the sounds were randomly scheduled and placed in different 3D locations to avoid learning effects on sound profiles across trials. In Unity, RW sounds were spatialized based on their 3D locations, while VR sounds were rendered stereo, simulating the experience of using a hearing device. For Prioritizer, the known sound schedules in Unity allowed us to adjust sound characteristics (e.g., volume, pitch, and duration.) to differentiate overlapping sounds or reschedule them to avoid conflict.
For Ambience Builder, objects producing sound were labeled as either RW or VR, with noise cancellation effects applied only to RW objects. We also used a headphone-blocking volume level $\eta$ to simulate the volume level reduced when noise-canceling headphones block the ears. By adjusting the level of transparency $\tau$ and the headphone-blocking volume level $\eta$, the volume of RW sound sources $S_{\text{new}}$ was determined by:
@@ 5.3. Technical Details in Unity Implementation
where we set the default volume of sounds $S_{\text{default}}=0.5$ to allow room for further manipulations (e.g., increasing volume by Feature Shifter), as in Unity the volume ranges from 0 to maximum 1. For the headphone-blocking volume-reduced level $\eta$, we set $\eta=0.75$ in our implementation, calculated based on the maximum amount of noise that can be canceled by wearing active noise-cancellation (ANC) headphones (45 dB, as reported by (Solutions, 2023)), and the average noise level in real-world environments (60 dB, as reported in (for Disease Control and Prevention, 2023)). Furthermore, a high pass filter is added to all the RW sound sources and its frequency cutoff $C$ can be determined based on the transparency level $\tau$:
@@ 5.4. Apparatus
The studies were conducted in person, where participants used Apple AirPods Max provided by us (see Figure 5). We allowed participants to choose whether to use the AirPods Max’s active noise cancellation mode, as some BVI individuals in our pilot study found it uncomfortable. However, we avoided the transparency mode to prevent the amplified real-world noises from distracting the participants. The studies took place in a soundproof, enclosed space to minimize distractions from external sounds. Participants completed the study tasks using a wireless keyboard, which, along with the AirPods Max, was connected via Bluetooth to a smartphone or tablet running our Unity application.
@@
In our study, participants wore headphones, engaged in pre-defined scenarios, and pressed specific keys upon hearing corresponding sounds.
@@ 5.5. Tasks and Procedure
After being welcomed and presented with our study’s informed consent and procedure, participants experienced three simulated scenarios (RW-Focused, VR-Focused, Fully-Mixed) under three conditions: Full Transparency (FT), Noise Cancellation (NC), and Sound Unblending (SU). The order of scenarios and conditions was counterbalanced for the 18 participants. Participants repeated each condition for five trials, which constituted a session.
@@ 5.7. Results
With Sound Unblending, participants achieved a significantly higher success rate and lower time delay in identifying sounds than FT and NC, but sound manipulations may confuse users’ interpretation of the sound content.
Overall, we found that Condition had a significant main effect on both Success Rate ($F$(2,16700)=327.58, $p$¡0.0001) and Delay Time ($F$(2,13242)=88.02, $p$¡0.0001) in the three scenarios (Figure 65(a)). Post-hoc Tukey’s pairwise test revealed that SU (M=0.897) resulted in a significantly higher Success Rate than FT (M=0.770) and NC (M=0.713), and FT was also significantly higher than NC, $p$¡0.0001; for Delay Time, SU (M=1.43s) was significantly lower than FT (M=1.66s) and NC (M=1.72s), and FT was significantly lower than NC, $p$¡0.0001. It was expected that SU would increase sound awareness in both RW and VR, and all participants commended the benefits of Sound Unblending: “SU achieves proper volume without the problem of sounds being too noisy or mixed up” (B16, RW-Focused), “SU is quite authentic and echos with my custom … I like to wear headphones on the right ear only when walking” (B12, RW-Focused), “I like SU as I also used to listen to certain things on one side” (B8, VR-Focused), and “I like it since it makes sounds very apparent, like notification of table cleaning, and the real people’s voices are also different from broadcast” (B3, Fully-Mixed). Also, the reason that FT led to better performance than NC was due to the nature of NC that blocks out RW sounds, as remarked by most participants and evidenced by the Success Rate of RW (M=0.693) and VR (M=0.895) sounds, and Delay Time of RW (M=1.81s) and VR (M=1.51s) sounds. Also, FT was preferable as it retained RW awareness and was more familiar to BVI participants, echoing the results in section 3.2.2 that BVI people emphasized the importance of retaining RW awareness when surveying different headphones.
However, participants also commented that Sound Unblending may confuse their perception and interpretation of RW sounds, such as “… the post-processed gentle drilling made me hard to tell the scale of the construction sites” (B10, RW-Focused), “It is weird … knocking sounds are far away from me” (B8,VR-Focused), and “The robot-like filter is distinguishable but distorts the content … I need to pay a lot of attention” (B12, Fully-Mixed). This highlights a trade-off between enhancing sound awareness and preserving the fidelity of certain sound characteristics when manipulating sounds.
@@
We found that Scenario had a significant main effect on both Success Rate ($F$(2,16700)=83.22, $p$¡0.0001) and Time Delay ($F$(2,13242)=810.83, $p$¡0.0001). Post-hoc Tukey’s pairwise test revealed that participants had significantly higher Success Rate in RW-Focused (M=0.848) than in VR-Focused (M=0.756) and Fully-Mixed scenarios (M=0.778), $p$¡0.0001 (Figure 65(b)). Furthermore, Fully-Mixed was significantly higher than VR-Focused, $p$=0.001. Participants also had significantly lower Delay Time in RW-Focused (M=1.24s) than in VR-Focused (M=1.36s) and Fully-Mixed (M=2.14s) scenarios, $p$¡0.0001; VR-Focused was significantly lower than Fully-Mixed, $p$¡0.0001.
@@
Results of NASA-TLX with error bars showing 95% confidence intervals. ***=$p$¡0.0001, **=$p$¡0.001 and *=$p$¡0.01.
@@ 5.7.3. RQ3: How do the conditions affect participants’ performance differently across the scenarios?
In all scenarios, SU was more effective and preferable than FT and NC, but still, many participants had varied preferences on sound awareness and found the sound clarity and naturalness of FT and the quietness of NC useful.
@@
We also found a significant interaction effect (Figure 76(b)) between Condition and Scenario for both Success Rate ($F$(4,16700)=56.7, $p$¡0.0001) and Delay Time ($F$(4,13242)=23.96, $p$¡0.0001). Specifically, in the RW-Focused scenario, post-hoc Tukey’s pairwise test revealed that participants had significantly higher Success Rate with SU (M=0.926) than FT (M=0.861) and NC (M=0.757), and FT than NC, $p$¡0.0001 (Figure 76(a)); they also had significantly lower Delay Time with SU (M=1.18s) than NC (M=1.30s), $p$¡0.0001. This was evidenced by most participants (N=14) that SU was preferable and provided a clear sound awareness of both RW and VR sounds than FT and NC, as B18 stated “I love the first one [SU], because the ambient noise in the second one [FT] is excessively noisy, and the construction sound in the third one [NC] is almost inaudible, which is very dangerous.” On the other hand, four participants liked FT the most as “… the sound of the white cane is more distinctive. The drilling is more vivid, which makes me more aware of the situation on the road … It [FT] is helpful because the sound is pretty spatial, I am able to perceive the composition of the environment” (B14). Participants also indicated that while NC made instructions clearer, it also lost information from RW, which was unfavorable in navigation scenario.
@@
In the VR-Focused scenario, post-hoc Tukey’s pairwise test revealed that participants achieved a significantly higher Success Rate with SU (M=0.874) than FT (M=0.779) and NC (M=0.615), and FT than NC, $p$¡0.0001 (Figure 76(a)); they also had significantly lower Delay Time with SU (M=1.26s) than FT (M=1.42s, $p$=0.0065) and NC (M=1.44s, $p$¡0.0001). Fourteen participants preferred and performed better in SU for its distinct audio channels, with the audio handbook on the left and the supervisor’s voice note on the right: “All details can be observed. The separate audio channel design is quite good, especially in such a static environment … it allowed me to distinguish different sounds from different channels … ease my burden” (B14). Another three participants liked SU and FT equally as FT retained the fidelity of sounds, as remarked by B8: “Both SU and FT have aspects I like. FT makes it a real-life scenario, while the information in SU is very clear.” One participant (B13) preferred NC for its quieter environment due to his high sensitivity to sounds.
@@
In the Fully-Mixed scenario, post-hoc Tukey’s pairwise test revealed that participants performed significantly better with SU (M=0.891) than FT (M=0.68) and NC (M=0.762), and NC than FT, $p$¡0.0001 (Figure 76(a)); they also had significantly lower Delay Time with SU (M=1.82s) than FT (M=2.37s, $p$¡0.0001) and NC (M=2.32s, $p$¡0.0001). Many participants (N=11) preferred SU for its sound clarity, as commented by B15: “It’s helpful and allows me to easily differentiate between broadcasts and human voices. Also, the sounds don’t seem to overlap as much.” However, some participants felt more pressured and overwhelmed in SU and preferred FT (N=5), as stated by B10: “I liked FT which has natural voices, and it has no earcons that might drown out human voices”, while others preferred NC (N=2), like B12: “The sound in this condition [NC] is clear and comes gradually, though it doesn’t make every sound very clear. The second one [SU] makes sounds too clear … It might cause me to overlook other important sounds.” In this complex scenario, participants’ preferences varied, and the enhanced clarity of specific sounds in SU raised concerns about potentially missing other crucial audio information.
@@
Sound Unblending overall reduced the cognitive load.
We found a significant main effect of Condition on the overall workload in NASA-TLX ($F$(2,136)=17.01, $p$¡0.0001). Post-hoc Tukey’s pairwise test revealed that SU (M=34.8) resulted in a significantly lower workload than both FT (M=45.4) and NC (M=48.2), $p$¡0.0001, but there was no significant difference between FT and NC ($p$=0.4637). The reason behind this result might be explained by the previous findings, where SU increased sound awareness via different manipulations; on the other hand, both FT and NC resulted in similar cognitive load, likely due to their inherent limitations, such as the multiple overlapping sounds in FT and the nearly-blocked RW sounds in NC. Interestingly, we found no significant differences between SU and the other two conditions for the overall cognitive load in the Fully-Mixed scenario.
This might be due to the result in section 5.7.3 that the stylized voices in SU sometimes made content unclear, or the enhanced clarity of specific sounds in SU might overshadow other crucial audio information.
5.7.5. RQ5: How are participants’ experiences and ways to further customize their soundscape for each scenario?
@@
First, most participants favored the natural presentation of RW sounds to maintain their perception of the real world, (Section 5.7.1). Yet, some participants desired more intelligent sound manipulations based on the RW context. For instance, B1 said in the RW-Focused scenario “I would only mute the voices of strangers but not all people as I want to hear the voices of someone I know”, and B10 also mentioned “I think earcon should be played only when the construction site is large enough, as typically, a small scale of construction would not be so noisy and discomforting.” Similarly, most participants also preferred keeping the voices of virtual people in the Fully-Mixed scenario unaltered, not imposing another sound filter on them. Instead, they may use other methods, such as echolocation or discerning voice nuances, to distinguish between RW and virtual reality (VR) sources.
@@
6. Example Applications with Sound Unblending
Based on insights from our studies, we developed three proof-of-concept prototypes to showcase the practicality and generalizability of Sound Unblending manipulations in different MR experiences. Though BVI people did not evaluate these applications, we hoped these applications stimulate further discussions and advancements in sound manipulation for future mixed-reality applications to promote accessibility. Note that these prototypes are functional. We highlight user interactions rather than technical specifics in this section. Please refer to our Video Figure for demonstrations.

### v2

@@ 5.1. Participants
Through word-of-mouth and public recruitment posts, we recruited 18 BVI participants (10 M and 8 F), aged 20 to 41 (mean=29.0), with a deep experience of using sounds in their lives. All had professional Orientation and Mobility (O&M) training and various sound-related experiences in other domains, like playing instruments, being voice actors, participating in orchestras, and broadcasting. Twelve were blind since birth, while six lost their vision later in life (Table 1). We refer to our BVI participants as B1-B18 in the following sections.
Table 1. Participant demographics information. O&M refers to Orientation and Mobility.
@@
B16 29 Female Blind, since birth. Light perception. O&M, Instrument, Making audio descriptions
@@ 5.2. Simulated Scenarios and Sound Manipulations
To fully explore the space of MR, we present three scenarios with different focuses on sound awareness: RW-Focused, VR-Focused, and Fully-Mixed. We simulated the three scenarios in Unity and used its built-in audio functions to prototype the six manipulators. In each scenario, there were four types of sounds, and participants were asked to complete sound identification tasks by pressing down specific keys on the keyboard upon hearing corresponding sounds. Participants were instructed to engage in the scenarios by prioritizing certain types of sounds relevant to the scenario’s objectives, such as concentrating on RW sounds in the RW-Focused scenario, or both RW and VR sounds in the Fully-Mixed scenario.
@@
[…]
Figure 2a A man wearing a black jacket is navigating the street with the white cane. Figure 2b The phone held by the man’s hands showing the sounds were spatialized on the mobile app. Figure 2c The first-person view of the man, including the road, street light, buildings and sidewalks. Figure 2d There are cars and crowds on the street and vendors on the sidewalk. Figure 2e There are construction sites between the road and the sidewalk. Figure 2f A close look at the manhole.
@@
  • Transparency Shift: applies half acoustic transparency (Figure 1d.1).
  • Envelope Shift: prioritizes the four sound types by volume: white cane on the manhole, drilling, navigation instructions, and ringtone (Figure 1d.2).
  • Position Shift: places navigation instructions on the left and ringtones on the right (Figure 1d.5).
  • Style Shift: applies low pass filters to drilling noises to be audibly comfortable (Figure 1d.3).
  • Time Shift: delays virtual sounds until after the RW sounds end (Figure 1d.4).
  • Sound Append: appends a short earcon when getting close to a construction site (Figure 1d.3).
@@
[…]
Figure 3 A global view of the Unity scene of VR-Focused scenario. Figure 3a A third-person view looking at the back of the person, who wearing headphones sitting in front of the monitor. Figure 3b A monitor showing some app icons and a notification popup window is on the top right corner of the screen. Figure 3c The space in front of the help desk. There are crowds of people and there is a speaker on the wall and cabinets and round tables in the sides or the corners of the room. Figure 3d A person come to the help desk and knock on the edge of the table Figure 3e There is text showing “Announcement: A reminder that our event is happening at…”. This comes from the speaker on the wall.
@@
  • Transparency Shift: applies full noise cancellation by default and dynamically switches to half acoustic transparency during public announcements.
  • Envelope Shift: prioritizes the four sound types by volume: knocking, public announcements, audio handbook, and voice notes.
  • Position Shift: places audio handbook on the left and voice notes on the right.
  • Time Shift: delays virtual sounds until after the RW sounds end.
@@
[…]
Figure 4 A global view of the Unity scene of Fully-Mixed scenario. Figure 4a A third-person view looking at the back of the person, who wears headphones sitting in front of the round table with people. There are text showing “Virtual Broadcast” and “Virtual Speaker”. Figure 4b A zoom out view of the reception room. There are crowds of people, including virtual and physical speakers. The virtual ones are presented by holo presence. Figure 4c Some speakers sit on the chair while others stand up and talk. Figure 4d In the round table, there are also virtual and physical attendees, there are foods on the table and glasses of water. Figure 4e There are waitstaff coming to the table with the food cart.
@@
  • Transparency Shift: applies half acoustic transparency.
  • Envelope Shift: prioritizes the four sound types using volume in the order of virtual/real people’s voices, table cleaning, and virtual broadcasts.
  • Position Shift: places virtual broadcasts on the right.
  • Style Shift: applies a heavier old-school telephone effect to virtual voices.
  • Time Shift: delays virtual sounds until after the RW sounds end.
  • Sound Append: appends two separate earcons for table cleaning and virtual broadcasts.
5.3. Technical Details on the Unity Implementation
In each trial, the sounds were randomly scheduled and placed in different 3D locations to avoid learning effects on sound profiles across trials. In Unity, RW sounds were spatialized based on their 3D locations. VR sounds were also able to be rendered spatially but mainly on either the left or right ear, or both in our study (Section 5.2), using commodity hearing devices. For Time Shift, the known sound schedules in Unity allowed us to adjust sound characteristics (e.g., volume, pitch, and duration) to differentiate overlapping sounds or reschedule them to avoid conflict.
For Transparency Shift, objects producing sound were labeled as either RW or VR, with noise cancellation effects applied only to RW objects. We also used a headphone-blocking volume level $\eta$ to simulate the volume level reduced when noise-canceling headphones block the ears. By adjusting the level of transparency $\tau$ and the headphone-blocking volume level $\eta$, the volume of RW sound sources $S_{\text{new}}$ was determined by:
@@ 5.3. Technical Details on the Unity Implementation
where we set the default volume of sounds $S_{\text{default}}=0.5$ to allow room for further manipulations (e.g., increasing volume by Envelope Shift), as in Unity the volume ranges from 0 to maximum 1. For the headphone-blocking volume-reduced level $\eta$, we set $\eta=0.75$ in our implementation, calculated based on the maximum amount of noise that can be canceled by wearing active noise-cancellation (ANC) headphones (45 dB, as reported by (Solutions, 2023)), and the average noise level in real-world environments (60 dB, as reported in (for Disease Control and Prevention, 2023)). Furthermore, a high pass filter is added to all the RW sound sources and its frequency cutoff $C$ can be determined based on the transparency level $\tau$:
@@ 5.4. Apparatus
The studies were conducted in person, where participants used Apple AirPods Max provided by us (see Figure 5). We allowed participants to choose whether to use the AirPods Max’s active noise cancellation mode, as some BVI individuals in our pilot study found it uncomfortable. However, we avoided the transparency mode to prevent the amplified real-world noises from distracting the participants. The studies took place in a soundproof, enclosed space to minimize distractions from external sounds. Participants completed the study tasks using a wireless keyboard, which, along with the AirPods Max, was connected via Bluetooth to a smartphone or tablet running our Unity application. Users’ head movements were tracked using the gyroscope data from the AirPods Max to enable spatial audio.
@@
[…]
Figure 5 A user sitting in front of the table. There are a tablet and keyboard on the table. The RW-Focused scenario is shown on the tablet. There is text in the figure “Imaging you are walking down the street, and there are several sounds. Press down keys …. When you hear …”
@@
[…]
Figure 6a This chart shows the results by conditions. Detailed results can be referred to Section 5.1. Figure 6b This chart shows the results by using the scenarios. Detailed results can be referred to Section 5.2.
@@ 5.5. Tasks and Procedure
After being welcomed and presented with our study’s informed consent and procedure, participants experienced three simulated scenarios (RW-Focused, VR-Focused, Fully-Mixed) under three conditions: Full Transparency (FT), Noise Cancellation (NC), and SoundShift (SS). The order of scenarios and conditions was counterbalanced for the 18 participants. Participants repeated each condition for five trials, which constituted a session.
@@ 5.6. Dependent Measures and Data Analysis
Figure 7. Overall results for RQ3 based on Condition and Scenario, including (a) Success Rate (top, the higher the better) and Delay Time (bottom, the lower the better) with error bars showing 95% confidence intervals. ***=$p$¡0.0001. (b) Interaction effects on Success Rate and Delay Time between Condition and Scenario.
@@
Figure 7 This chart shows the results by examining both conditions and scenarios. Detailed results can be referred to Section 5.3. (a) Success Rate (top, the higher the better) and Delay Time (bottom, the lower the better) with error bars showing 95% confidence intervals. ***=$p$¡0.0001. (b) Interaction effects on Success Rate and Delay Time between Condition and Scenario.
@@ 5.7. Results
With SoundShift, participants achieved a significantly higher success rate and lower time delay in identifying sounds than FT and NC, but sound manipulations may confuse users’ interpretation of the sound content.
Overall, we found that Condition had a significant main effect on both Success Rate ($F$(2,16700)=327.58, $p$¡0.0001) and Delay Time ($F$(2,13242)=88.02, $p$¡0.0001) in the three scenarios (Figure 6a). Post-hoc Tukey’s pairwise test revealed that SS (M=0.897) resulted in a significantly higher Success Rate than FT (M=0.770) and NC (M=0.713), and FT was also significantly higher than NC, $p$¡0.0001; for Delay Time, SS (M=1.43s) was significantly lower than FT (M=1.66s) and NC (M=1.72s), and FT was significantly lower than NC, $p$¡0.0001. It was expected that SS would increase sound awareness in both RW and VR, and all participants commended the benefits of SoundShift: “SS achieves proper volume without the problem of sounds being too noisy or mixed up” (B16, RW-Focused), “SS is quite authentic and echos with my custom … I like to wear headphones on the right ear only when walking” (B12, RW-Focused), “I like SS as I also used to listen to certain things on one side” (B8, VR-Focused), and “I like it since it makes sounds very apparent, like notification of table cleaning, and the real people’s voices are also different from broadcast” (B3, Fully-Mixed). Also, the reason that FT led to better performance than NC was due to the nature of NC that blocks out RW sounds, as remarked by most participants and evidenced by the Success Rate of RW (M=0.693) and VR (M=0.895) sounds, and Delay Time of RW (M=1.81s) and VR (M=1.51s) sounds. Also, FT was preferable as it retained RW awareness and was more familiar to BVI participants, echoing the results in section 3.2.2 that BVI people emphasized the importance of retaining RW awareness when surveying different headphones.
However, participants also commented that SoundShift may confuse their perception and interpretation of RW sounds, such as “… the post-processed gentle drilling made me hard to tell the scale of the construction sites” (B10, RW-Focused), “It is weird … knocking sounds are far away from me” (B8,VR-Focused), and “The robot-like filter is distinguishable but distorts the content … I need to pay a lot of attention” (B12, Fully-Mixed). This highlights a trade-off between enhancing sound awareness and preserving the fidelity of certain sound characteristics when manipulating sounds.
@@
We found that Scenario had a significant main effect on both Success Rate ($F$(2,16700)=83.22, $p$¡0.0001) and Time Delay ($F$(2,13242)=810.83, $p$¡0.0001). Post-hoc Tukey’s pairwise test revealed that participants had significantly higher Success Rate in RW-Focused (M=0.848) than in VR-Focused (M=0.756) and Fully-Mixed scenarios (M=0.778), $p$¡0.0001 (Figure 6b). Furthermore, Fully-Mixed was significantly higher than VR-Focused, $p$=0.001. Participants also had significantly lower Delay Time in RW-Focused (M=1.24s) than in VR-Focused (M=1.36s) and Fully-Mixed (M=2.14s) scenarios, $p$¡0.0001; VR-Focused was significantly lower than Fully-Mixed, $p$¡0.0001.
@@
[…]
Figure 8 This chart shows NASA-TLX results for RQ4. Detailed results can be referred to in Section 5.4.
@@ 5.7.3. RQ3: How do the conditions affect participants’ performance differently across the scenarios?
In all scenarios, SS was more effective and preferable than FT and NC, but still, many participants had varied preferences on sound awareness and found the sound clarity and naturalness of FT and the quietness of NC useful.
@@
We also found a significant interaction effect (Figure 7b) between Condition and Scenario for both Success Rate ($F$(4,16700)=56.7, $p$¡0.0001) and Delay Time ($F$(4,13242)=23.96, $p$¡0.0001). Specifically, in the RW-Focused scenario, post-hoc Tukey’s pairwise test revealed that participants had significantly higher Success Rate with SS (M=0.926) than FT (M=0.861) and NC (M=0.757), and FT than NC, $p$¡0.0001 (Figure 7a); they also had significantly lower Delay Time with SS (M=1.18s) than NC (M=1.30s), $p$¡0.0001. This was evidenced by most participants (N=14) that SS was preferable and provided a clear sound awareness of both RW and VR sounds than FT and NC, as B18 stated “I love the first one [SS], because the ambient noise in the second one [FT] is excessively noisy, and the construction sound in the third one [NC] is almost inaudible, which is very dangerous.” On the other hand, four participants liked FT the most as “… the sound of the white cane is more distinctive. The drilling is more vivid, which makes me more aware of the situation on the road … It [FT] is helpful because the sound is pretty spatial, I am able to perceive the composition of the environment” (B14). Participants also indicated that while NC made instructions clearer, it also lost information from RW, which was unfavorable in navigation scenario.
@@
In the VR-Focused scenario, post-hoc Tukey’s pairwise test revealed that participants achieved a significantly higher Success Rate with SS (M=0.874) than FT (M=0.779) and NC (M=0.615), and FT than NC, $p$¡0.0001 (Figure 7a); they also had significantly lower Delay Time with SS (M=1.26s) than FT (M=1.42s, $p$=0.0065) and NC (M=1.44s, $p$¡0.0001). Fourteen participants preferred and performed better in SS for its distinct audio channels, with the audio handbook on the left and the supervisor’s voice note on the right: “All details can be observed. The separate audio channel design is quite good, especially in such a static environment … it allowed me to distinguish different sounds from different channels … ease my burden” (B14). Another three participants liked SS and FT equally as FT retained the fidelity of sounds, as remarked by B8: “Both SS and FT have aspects I like. FT makes it a real-life scenario, while the information in SS is very clear.” One participant (B13) preferred NC for its quieter environment due to his high sensitivity to sounds.
@@
In the Fully-Mixed scenario, post-hoc Tukey’s pairwise test revealed that participants performed significantly better with SS (M=0.891) than FT (M=0.68) and NC (M=0.762), and NC than FT, $p$¡0.0001 (Figure 7a); they also had significantly lower Delay Time with SS (M=1.82s) than FT (M=2.37s, $p$¡0.0001) and NC (M=2.32s, $p$¡0.0001). Many participants (N=11) preferred SS for its sound clarity, as commented by B15: “It’s helpful and allows me to easily differentiate between broadcasts and human voices. Also, the sounds don’t seem to overlap as much.” However, some participants felt more pressured and overwhelmed in SS and preferred FT (N=5), as stated by B10: “I liked FT which has natural voices, and it has no earcons that might drown out human voices”, while others preferred NC (N=2), like B12: “The sound in this condition [NC] is clear and comes gradually, though it doesn’t make every sound very clear. The second one [SS] makes sounds too clear … It might cause me to overlook other important sounds.” In this complex scenario, participants’ preferences varied, and the enhanced clarity of specific sounds in SS raised concerns about potentially missing other crucial audio information.
@@
SoundShift overall reduced the cognitive load.
We found a significant main effect of Condition on the overall workload in NASA-TLX ($F$(2,136)=17.01, $p$¡0.0001). Post-hoc Tukey’s pairwise test revealed that SS (M=34.8) resulted in a significantly lower workload than both FT (M=45.4) and NC (M=48.2), $p$¡0.0001, but there was no significant difference between FT and NC ($p$=0.4637). The reason behind this result might be explained by the previous findings, where SS increased sound awareness via different manipulations; on the other hand, both FT and NC resulted in similar cognitive load, likely due to their inherent limitations, such as the multiple overlapping sounds in FT and the nearly-blocked RW sounds in NC. Interestingly, we found no significant differences between SS and the other two conditions for the overall cognitive load in the Fully-Mixed scenario.
This might be due to the result in section 5.7.3 that the stylized voices in SS sometimes made content unclear, or the enhanced clarity of specific sounds in SS might overshadow other crucial audio information.
5.7.5. RQ5: How do participants describe their experiences and ways to further customize their soundscape for each scenario?
@@
First, most participants favored the natural presentation of RW sounds to maintain their perception of the real world, (Section 5.7.1). Yet, some participants desired more intelligent sound manipulations based on the RW context. For instance, B1 said in the RW-Focused scenario “I would only mute the voices of strangers but not all people as I want to hear the voices of someone I know”, and B10 also mentioned “I think earcon should be played only when the construction site is large enough, as typically, a small scale of construction would not be so noisy and discomforting.” Similarly, most participants also preferred keeping the voices of virtual people in the Fully-Mixed scenario unaltered, not imposing another sound filter on them. Instead, they may use other methods, such as sound localization or discerning voice nuances, to distinguish between RW and virtual reality (VR) sources.
@@
6. Example Applications with SoundShift
Based on insights from our studies, we developed three proof-of-concept prototypes to showcase the practicality and generalizability of SoundShift manipulations in different MR experiences. Though BVI people did not evaluate these applications, we hoped these applications stimulate further discussions and advancements in sound manipulation for future mixed-reality applications to promote accessibility. Note that these prototypes are functional. We highlight user interactions rather than technical specifics in this section. Please refer to our Video Figure for demonstrations.